The probability of making a 6-ft putt is 60%. How many times should you expect to make it if you putt the flowing numbers?

1.) 5times
2.) 10 times
3.) 30 times
4.) 160 times






Thanks so much

Answer by vleith(2517)

If the probability is 60%, that says 60 out of every 100 putts goes in.
Just multiply the number of putts by 60/100 (0.6)
5*0.6 = 3
10*0.6 = 6
30 * 0.6 = 18
You do the last one